OpenStudy (anonymous):

.Find all solutions to the equation. 7 sin2x - 14 sin x + 2 = -5

OpenStudy (campbell_st):

rewrite the equation. It is basically a quadratic equation \[7\sin^2(x) - 14\sin(x) + 7 = 0\] next divide every term by 7 \[\sin^2(x) - 2\sin(x) + 1 = 0\] you now have a perfect square which can solved by factoring